i) No change in either equilibrium price or equilibrium quantity nsufficient information to determine the effect on price or quantity hich of the following could explain a decrease in the price and quantity sold of gasoline over a summer long weekend? a) A decrease in Demand effect due to weather greater than the effect of a decrease b) A decrease in Demand effect due to weather less than the effect of a decrease c) An increase in Demand effect due to weather greater than the effect of a decrease d) An increase in Demand effect due to weather less than the effect of a decrease e) None of the above Questions 14 to 19 refer to the Demand and Supply functions below for Ontario Eggs per in the price of oil in the price of oil in the price of oil in the price of oil dozen in Toronto per day. P 15-0.004Q and P- 2+0.001Q 14. What is the competitive equilibrium quantity sold of Eggs (dozen) per day?
Answer #1

Answer : 13) The answer is option a.

Because if demand decrease too much then the market face a situation of excess supply. For this excess supply the price level decrease. If decrease in quantity demanded is higher than the decrease in price level then the market of gasoline face a situation of too much excess supply. As a result, the price level of gasoline decrease more. Therefore, option a is correct.

14) In competitive market at equilibrium, Demand = Supply.

=> 15 - 0.004Q = 2 + 0.001Q

=> 15 - 2 = 0.001Q + 0.004Q

=> 13 = 0.005Q

=> Q = 13 / 0.005

=> Q = 2,600

Therefore, the sold equilibrium quantity of competitive market is 2,600 Eggs (dozen) per day.